OpenStudy (anonymous):

elementry linear equation: using De Moivre's theorem,express cos4 theter and sin theter in terms of sines and cosines of multiples of theter

OpenStudy (dumbcow):

\[\cos(4\theta)+i \sin(4\theta) = (\cos \theta+i \sin \theta)^{4}\] is this what you mean?

OpenStudy (anonymous):

@Lesco , is it?

OpenStudy (anonymous):

yes, thank you so much.how can i determine the sixth roots of 8 as well as of -8.

OpenStudy (dumbcow):

\[x = \sqrt[6]{\pm 8}\] \[x^{6} \pm 8 = 0\] Factor using sum/difference of cubes \[(x^{2} \pm 2)(x^{4} \mp 2x^{2}+4)=0\] From here you can solve for 2 roots from 1st term for 2nd term, use quadratic formula

OpenStudy (dumbcow):

i think you can factor that further but i can't think of it at the moment

OpenStudy (anonymous):

thank you very much dumbcow, its been hard for to complete these questions, I'm new in linear equations ,i need help ..can you assist me with my assignment?

OpenStudy (dumbcow):

yes, i can help with questions you post